OpenStudy (anonymous):

9a-b-2a-10b A)-7a+11b B)11a+9b C)-11a-9b D)7a-11b Help please

OpenStudy (texaschic101):

All you have to do here is combine like terms. Do you know what like terms are ?

OpenStudy (anonymous):

no

OpenStudy (texaschic101):

Like terms are terms that have EXACTLY the same letters. Example : 2x and 3x are like terms and can be added to equal 5x 3x and 4y are not like terms and you cannot do anything with them. So lets look at our problem... 9a - b - 2a - 10b can you find the like terms ?

OpenStudy (anonymous):

9a-2a=7a?

OpenStudy (texaschic101):

correct.....9a - 2a = 7a are there any more like terms ?

OpenStudy (anonymous):

I don't think so

OpenStudy (texaschic101):

what about -b - 10b ? Keeping in mind that -b is the same as -1b.

OpenStudy (anonymous):

so D?

OpenStudy (texaschic101):

yes...because.. 9a - 2a = 7a -1b - 10b = -11b put it all together and you get : 7a - 11b

OpenStudy (anonymous):

thank you so much

OpenStudy (texaschic101):

anytime :)